OpenStudy (anonymous):

any help An airplane flew 4 hours with a 25 mph tail wind. The return trip against the same wind took 5 hours. Find the speed of the airplane in still air.

OpenStudy (btaylor):

Make a table: since d=d, we can set 4(r+25)=5(r-25) 4r+100 = 5r-125 r=225 mph